# Reptile trapezoids

### v1

@@ 1 Introduction
A geometric figure is said to be rep-$n$ if it can be dissected into $n$ similar copies congruent to each other. When a figure is rep-$n$ for some $n\geq 2$, it is called a reptile. In [1], C. D. Langford proposed the problem of characterizing reptile polygons while giving some examples of them, three of which are reptile trapezoids R1, R2, and R3 illustrated in Figure 1. Later on, S. W. Golomb presented further examples of reptile polygons in [2], including the ‘sphinx’, so far the only known non-convex reptile polygon with an odd number of sides according to [3, Kapitel 5]. Furthermore, in [3, Section 3.2], I. Osburg provided a method for constructing a non-convex reptile $2n$-gon with $n\geq 3$, showing that there are infinitely many non-convex reptile polygons.
On the other hand, every convex reptile polygon is known to be either a triangle or a quadrilateral according to the results of U. Betke in [4] (as cited in [5]) and of Osburg in [3, Satz 2.23]. Osburg also provided a narrower characterization for reptile quadrilaterals in [3, Satz 2.9 and Folgerung 3.2]: a quadrilateral is a reptile only if it is convex and is either a trapezoid or a cyclic quadrilateral. In addition, as proven by M. Laczkovich in [5, Corollary 1.3], it turned out that only trapezoids can be a reptile quadrilateral. Consequently, the only problem for classifying convex reptile polygons is finding all reptile trapezoids. I could not find the exact reference, but except for parallelograms, there seem to be only four known examples of reptile trapezoids so far, as illustrated in Figure 1, where the last one, R4, was mentioned by L. Sallows in [6].
@@
The structure of this paper is as follows. In Section 2, we introduce some terminologies and notations that will be frequently used throughout the discussion. Then, in Section 3, we describe the main results of this paper in advance, followed by a comment for further research and a proposal for some related conjectures. In Section 4, we discuss the properties and lemmas essential to the proof of the main results. Using them, we prove Theorem 3.1 in Section 5. We then utilize Theorem 3.1 and the lemmas in Section 4 to prove Theorem 3.2 in Section 6. Furthermore, in Section 7, we prove Corollary 3.4 using the result of Theorem 3.1 and prove Corollary 3.5 similarly in Section 8.1. Finally, we prove Theorem 3.6 throughout Section 8.2, 8.3 and 8.4, and prove Theorem 3.7 in Section 8.5.
@@ 3 Main results
The right trapezoid $R\left(\frac{\pi}{3},\frac{1}{8}\right)$ is rep-25 but not rep-$n$ for all $251<n<$ 25.
@@
The outline of the proof of Theorem 3.1, 3.2, and 3.6 is as follows. In essence, we prove Theorem 3.1 by contrapositive and proof by contradiction. Given a unit trapezoid $T$ not satisfying the necessary conditions in the theorem, we assume, for the sake of contradiction, that there is a tiling of $\mu T$ ($\mu\geq 3$) with congruent copies of $T$ (that is, $T$ is rep-$\mu^{2}$). We then show that the trapezoid $T$ or the given tiling satisfies some properties. By some related lemmas and corollaries, this implies the non-existence of tilings of $\mu T$ with congruent copies of $T$, which is a contradiction. This suffices to conclude that $T$ is not a reptile even though we neglected cases when $\mu$ equals $2$ or is the square root of some square-free integer instead of being an integer; note that if $T$ is rep-$n$, then it is also rep-$n^{m}$ for all integers $m\geq 2$ because, for each integer $i\geq 0$, the trapezoid $n^{(i+1)/2}T$ can be tiled with $n$ congruent copies of $n^{i/2}T$ if $T$ is rep-$n$. The other two theorems are proven straightforwardly using some related lemmas and corollaries, some of which are also used for proving Theorem 3.1.
@@ 7 Angular condition for reptile $I(\vartheta,a)$ with rational $\vartheta/\pi$
In this section, we prove Corollary 3.4. Suppose that $I(\vartheta,a)$ is a reptile. What immediately follows from Theorem 3.1 is that both $a$ and $a+2\cos{\vartheta}$ ) are rational, and thus so is $\cos{\vartheta}$ ). According to the result of P. Tangsupphathawat in [7, Theorem 3.3], if $\Theta$ is an acute angle such that both $\Theta/\pi$ and $\cos{\Theta}$ ) are rational, then $\Theta=\pi/3$. This shows that $\vartheta=\pi/3$ if $\vartheta/\pi\in\mathds{Q}$. This ends the proof of the corollary.
@@ 8.2 Angular condition for reptile $R(\vartheta,a)$ with rational $\vartheta/\pi$
In [7, Theorem 3.3], it is shown that if $\Theta$ is an acute angle such that $\Theta/\pi\in\mathds{Q}$ and $\cos\Theta$ is an algebraic integer of degree at most $2$, then $\Theta$ equals one of $\frac{\pi}{6},\frac{\pi}{5},\frac{\pi}{4},\frac{\pi}{3},$ and $\frac{2\pi}{5}$. Thus, if $R(\vartheta,a)$ is a reptile and $\vartheta/\pi$ is rational, then $\vartheta$ equals one of these five values. At first, we can rule out $\frac{\pi}{5}$ and $\frac{2\pi}{5}$ since the sine of each of these two values is not an algebraic integer of degree at most $2$ as the following equations show (in fact, both are algebraic integers of degree $4$).

### v2

# Jin Heo    Jin Heo

@@ 1 Introduction
A geometric figure is said to be rep-$n$ if it can be dissected into $n$ similar copies congruent to each other. When a figure is rep-$n$ for some $n\geq 2$, it is called a reptile. In [1], C. D. Langford proposed the problem of characterizing reptile polygons while giving some examples of them, three of which are reptile trapezoids R1, R2, and R3 illustrated in Figure 1. Later on, S. W. Golomb presented further examples of reptile polygons in [2], including the ‘sphinx,’ so far the only known non-convex reptile polygon with an odd number of sides according to [3, Kapitel 5]. Furthermore, in [3, Section 3.2], I. Osburg provided a method for constructing a non-convex reptile $2n$-gon with $n\geq 3$, showing that there are infinitely many non-convex reptile polygons.
On the other hand, every convex reptile polygon is known to be either a triangle or a quadrilateral according to the results of U. Betke in [4] (as cited in [5]) and of Osburg in [3, Satz 2.23]. Osburg also provided a narrower characterization for reptile quadrilaterals in [3, Satz 2.9 and Folgerung 3.2]: a quadrilateral is a reptile only if it is convex and is either a trapezoid or a cyclic quadrilateral. In addition, M. Laczkovich made a further breakthrough in [5, Corollary 1.3], proving that only trapezoids can be a reptile quadrilateral. Consequently, the only problem for classifying convex reptile polygons is finding all reptile trapezoids. I could not find the exact reference, but except for parallelograms, there seem to be only four known examples of reptile trapezoids so far, as illustrated in Figure 1, where the last one, R4, was mentioned by L. Sallows in [6].
@@
The structure of this paper is as follows. In Section 2, we introduce some terminologies and notations that will be frequently used throughout the discussion. Then, in Section 3, we describe the main results of this paper in advance, followed by a comment for further research and a proposal for some related conjectures. In Section 4, we discuss the properties and lemmas essential to the proof of the main results. Using them, we prove Theorem 3.1 in Section 5. We then utilize the results of Theorem 3.1 and the lemmas in Section 4 to prove Theorem 3.2 in Section 6. Furthermore, in Section 7, we prove Corollary 3.4 using the result of Theorem 3.1 and prove Corollary 3.5 similarly in Section 8.1. Finally, we prove Theorem 3.6 throughout Section 8.2, 8.3 and 8.4, and prove Theorem 3.7 in Section 8.5.
@@ 3 Main results
The right trapezoid $R\left(\frac{\pi}{3},\frac{1}{8}\right)$ is rep-25 but not rep-$n$ for all integers $251<n<$ 25.
@@
The outline of the proof of the main results is as follows. As a preliminary step, we first consider a polygon containing a specific polygonal line in its boundary and some area surrounded by the polygonal line in its region. Then, provided that the polygon is tiled with congruent copies of a given unit trapezoid $T$, we observe how the tiles should cover the polygonal line and a neighboring area if $T$ or the given tiling satisfies some properties; this is similar to the argument of Laczkovich in [5, Section 4] that concentrates on a local structure of a polygon. Based on this observation, we prove that such properties that are met by $T$ or the tiling lead to the non-existence of certain tilings of trapezoids in some families consisting of those obtained from a larger copy of $T$ by elongating or shortening its bases. During this step, we distinguish the case when $T$ is a right trapezoid from when it is not, for a right internal angle of $T$ necessitates a different approach.
After the preliminary step, we prove Theorem 3.1 by contradiction. For the sake of contradiction, we assume that congruent copies of a given unit trapezoid $T$ tile its larger copy $\mu T$ for some integer $\mu\geq 3$ (i.e. $T$ is rep-$\mu^{2}$), but at least one of $\left|\mathbf{ub}\left(T\right)\right|$ and $\left|\mathbf{lb}\left(T\right)\right|$ is not contained in $\mathds{Q}_{\left|\mathbf{sl}\left(T\right)\right|}$. We first derive a contradiction for the case when $\left|\mathbf{ub}\left(T\right)\right|\notin\mathds{Q}_{\left|\mathbf{sl}\left(T\right)\right|}$ and $\left|\mathbf{lb}\left(T\right)\right|\in\mathds{Q}_{\left|\mathbf{sl}\left(T\right)\right|}$ from a number-theoretical approach, which is similar to the arguments of S. L. Snover et al. in [7] and of Laczkovich in [5, Section 3]. For the remaining case, we use an analogous number-theoretical method to show that $T$ or the given tiling satisfies some of the properties introduced in the first step. We then derive a contradiction by applying the results in the preliminary step and, therefore, conclude that $T$ is not rep-$\mu^{2}$. This result is sufficient enough to imply that $T$ is not a reptile even though we neglected cases when $\mu$ equals either $2$ or the square root of some square-free integer; note that if $T$ is rep-$n$, then it is also rep-$n^{m}$ for all integers $m\geq 2$ because, for each integer $i\geq 0$, the trapezoid $n^{(i+1)/2}T$ can be tiled with $n$ congruent copies of $n^{i/2}T$ if $T$ is rep-$n$.
Based on the result of Theorem 3.1, we prove Corollary 3.5 by using a trigonometric equation. By additionally applying the result of P. Tangsupphathawat in [8, Theorem 3.3] to the main acute angle of a given unit trapezoid, we prove Corollary 3.4 and the angle-related part of Theorem 3.6.
The proof of Theorem 3.2 and the edge-related part of Theorem 3.6 is straightforward. We first assume that a given unit trapezoid $T$ is a reptile, particularly rep-$\mu^{2}$ for some integer $\mu\geq 3$. Similar to the proof of Theorem 3.1, we consider a tiling of a larger copy $\mu T$ with congruent copies of $T$. By applying the previous number-theoretical approach, we eliminate contradictory cases where some of the properties introduced in the first step are satisfied, giving us the desired results.
Finally, Theorem 3.7 will be proven by investigating one by one whether congruent copies of the unit trapezoid $T=R\left(\frac{\pi}{3},\frac{1}{8}\right)$ can tile its larger copy $\sqrt{n}T$ for some integer $251<n<$ 25.
@@ 7 Angular condition for reptile $I(\vartheta,a)$ with rational $\vartheta/\pi$
In this section, we prove Corollary 3.4. Suppose that $I(\vartheta,a)$ is a reptile. What immediately follows from Theorem 3.1 is that both $a$ and $a+2\cos{\vartheta}$ ) are rational, and thus so is $\cos{\vartheta}$ ). According to the result of Tangsupphathawat in [8, Theorem 3.3], if $\Theta$ is an acute angle such that both $\Theta/\pi$ and $\cos{\Theta}$ ) are rational, then $\Theta=\pi/3$. This shows that $\vartheta=\pi/3$ if $\vartheta/\pi\in\mathds{Q}$. This ends the proof of the corollary.
@@ 8.2 Angular condition for reptile $R(\vartheta,a)$ with rational $\vartheta/\pi$
In [8, Theorem 3.3], it is shown that if $\Theta$ is an acute angle such that $\Theta/\pi\in\mathds{Q}$ and $\cos\Theta$ is an algebraic integer of degree at most $2$, then $\Theta$ equals one of $\frac{\pi}{6},\frac{\pi}{5},\frac{\pi}{4},\frac{\pi}{3},$ and $\frac{2\pi}{5}$. Thus, if $R(\vartheta,a)$ is a reptile and $\vartheta/\pi$ is rational, then $\vartheta$ equals one of these five values. At first, we can rule out $\frac{\pi}{5}$ and $\frac{2\pi}{5}$ since the sine of each of these two values is not an algebraic integer of degree at most $2$ as the following equations show (in fact, both are algebraic integers of degree $4$).